OpenStudy (anonymous):

3x+8x=55 solve the equation

OpenStudy (anonymous):

11x = 55, x=5

OpenStudy (doc.brown):

\[3x = xxx\]\[8x = xxxxxxxx\]\[3x + 8x = xxx+xxxxxxxx\]\[xxxxxxxxxxx = 11x\]\[11x=55\]\[\frac{1}{11}\times 11x=55\times\frac{1}{11}\]\[\frac{11}{11}x=\frac{55}{11}\]\[1x=5\]\[x=5\]